# Interpretation of a Discrete de Rham method as a Finite Element System

### v1

@@ Abstract
We show that the DDR method can be interpreted as defining a computable consistent discrete $\mathrm{L}^{2}$ product on a conforming FES defined by PDEs. Without modifying the numerical method itself, this point of view provides an alternative approach to the analysis. The conformity and consistency properties we prove are stronger than those previously shown. We can also recover some of the other results that have been proved about DDR, from those that have already been proved, in principle, in the general context of FES. We also bring VEM, the Virtual Element Method, into the discussion.
@@ Introduction
We give a new presentation of the DDR method of [27, 9]. We interpret it as a numerical method based on discrete spaces defined piecewise by solving certain new Partial Differential Equations (PDEs), on each cell of a mesh. As such, the method is consistent in the sense of Strang’s lemmas, and conforming in the sense of producing subcomplexes of natural $\mathrm{L}^{2}$-based Hilbert complexes. Even though the solutions to these new local PDEs are not computable, the numerical method itself is ”fully discrete” (computable). This type of consistency and conformity are stronger than those previously proved for DDR. The interpretation also embeds the method in the Finite Element System framework, and allows to leverage the results already obtained in that setting.
@@
The framework of Finite Element Systems (FES) was introduced in [11] and expanded upon in [20, §5] (this independent section is available on arXiv [^1] S. H. Christiansen, Finite element systems of differential forms, arXiv:1006.4779, 2015.) and subsequent papers. It is an environment in which to define and analyze sequences of spaces of differential forms, that are finite elements in a broad sense. The underlying mesh can consist of cells (in particular polytopes) and the basis functions can be general (not necessarily based on polynomials). Thus one may call them polytopal discrete de Rham sequences, though this expression is given an a priori different meaning below. These spaces were conforming in the sense that the fields were $\mathrm{L}^{2}$-integrable with exterior derivative also $\mathrm{L}^{2}$-integrable, which enforces that traces, in the sence of pullbacks to faces, coincide (they are ”single-valued”). This encodes a type of partial continuity that is common for the mixed finite elements of electromagnetics and fluid flow [8], such as those of Raviart-Thomas-Nédélec. The latter were extended to differential forms in [32]. A synthesis is provided in Finite Element Exterior Calculus (FEEC) [3].
@@
Since then the framework of FES has been generalized to accomodate higher regularity of the fields [18] (as in so-called ”Stokes complexes” relevant to fluid flow) and other types of fields [19] (e.g. symmetric tensors and their associated curvature tensors, as in Riemannian geometry). The definition has a natural expression in the language of categories and a short summary is provided in [15]. It now covers all finite elements described according to Ciarlet’s definition, including non-conforming ones (such as Morley and Crouzeix-Raviart elements).
@@
For the purposes of this paper, however, the definition given in [11] is sufficient. The goal back then was to define a discrete de Rham finite element sequence on polytopal meshes that mimicked the lowest order mixed finite elements that correspond to Whitney forms. Indeed these were known only for simplices (and products of simplices handled by tensor product constructions). For this presentation will also refer to the more comprehensive presentation in [20, §5]. Short presentations can also be found in later publications, in particular [21, §2].
@@
The Virtual Element Method (VEM), reviewed in [5], has been developed with similar goals as FES and some similar basic principles, such as definition by recursive extension of the fields, from low dimensional cells. But, crucially, no submesh is used in VEM to define the discretization spaces or the numerical method. The challenge is to reconstruct enough information on the fields from the degrees of freedom, to obtain stable and consistent approximations of the bilinear forms defining the partial differential equation to solve. A connection between the mixed VEM (defined in dimension 3) of [4] and FES (in arbitrary dimension) was made in [16, §2.1] and will be used later.
@@
The close relationship between DDR and VEM has already been pointed out in [6]. In some versions these methods can be seen to have the same degrees of freedom, and they coincide with those of FES provided in [16, §2.1].
In this paper, we reinterpret DDR as a conforming FES equipped with computable consistent approximations of the $\mathrm{L}^{2}$ scalar products. The degrees of freedom of [16, §2.1] are used. This shows that these examples of DDR, VEM and FES are very similar, if not identical, in practice. They seem to differ mainly in terms of interpretation and analysis. The main novelty here is just that a new harmonic extension PDE, modifying the one used in [16, §2.1], is introduced to define the discrete spaces.
This point of view allows us to recover some of the results that have been proved for VEM or DDR [25, 26], as special cases of results already proved, in their principle, in the general context of FES. To sum up:
@@
  • Cohomological properties developed for FES, especially [20, Proposition 5.15], now carry over to DDR. This passage from local exactness (on each cell of the mesh), to globally correct cohomology (on the computational domain) illustrates that the existence of unisolvent degrees of freedom on a FES can be interpreted as a softness condition on sheaves (see [33, §III.2] for definitions).
  • Uniformly bounded commuting projections, building on [34, 10], have been described for FEEC [3, 23] and FES [20, §5.3]. They enable one to recover Poincaré-Friedrich estimates, which are basic to the stability analysis of mixed methods.
In the setting of FES one furthermore obtains refinements:
  • Translation estimates are given in [20, Proposition 5.71], and have applications to the equations of fluid flow.
  • Discrete Rellich compactness, as introduced by Kikuchi for Maxwell [7, §IV.19], holds generally, as proved first in [10, Corollary 5] with applications to the Hodge-Laplace eigenvalue problem on manifolds. See also [24] on this topic.
  • Discrete Sobolev injections hold [20, Propositions 5.69, 5.70] and have applications to, for instance, non-linear Maxwell wave equations as in [22]. See also [31] on this topic.
@@
Remark 0.1.
More recent constructions of local bounded commuting projections [29, 2] carry over to FES in a very natural way. The method of [29] extends to FES of differential forms and the method of [2] extends to general FES as in [19] (the latter method does not appeal to a wedge product). In fact the FES framework seems to identify exactly the algebraic properties of discretization spaces needed for these constructions. A general analysis of stability and compactness properties of these methods remains, in the context of FES, to be fleshed out. See [28] for the state of the art.
While we have little to add to the design of the DDR and VEM methods (a part from a new rationale for neglecting certain terms), we hope the FES point of view can contribute usefully to their analysis. We also emphasize that many of the technical computations that appear here without further ado, already appear exemplified in the above papers on DDR and VEM, though perhaps not in this order and generality.
@@ 1 Notations
The subspace of $X^{k}(T)$ consisting of forms with 0 traces on $\partial T$ is denoted $X^{k}_{0}(T)$. Subscript 0 will have the same meaning for other spaces as well. Elements on $X^{k}(T)$ have well defined traces, in negative index Sobolev spaces, only on codimension one faces $T^{\prime}\subseteq\partial T$. However, if these traces in turn are imposed to be in $X^{k}(T^{\prime})$, one can repeat the procedure and follow traces all the way down. One gets a space of forms with improved boundary regularity and some nice properties, such as [21, Proposition 2.4]. All our discrete spaces will have this type of extra boundary regularity.
The space of Whitney $k$-forms (which could also be attributed to Weil and de Rham) was originally presented in terms of partitions of unity, such as barycentric coordinates, but it can also be presented in terms of the Koszul operator [3, §3.2]. For each cell $T\in\mathcal{T}$, we let $b_{T}$ be its barycentre (we suppose it is ”computable”). On $T$, the Koszul operator $\kappa_{T}$ is the contraction by the vector field on $T$ given as $x\mapsto x-b_{T}$ (for $x\in T$). We adopt notations from [3]. The original lowest order Whitney forms are then $\mathcal{P}_{\!-}^{1}\Lambda^{k}(T)$. The high order version are denoted $\mathcal{P}_{\!-}^{r}\Lambda^{k}(T)$:
$\mathcal{P}_{\!-}^{r}\Lambda^{k}(T)=\{u\in\mathcal{P}^{r}\Lambda^{k}(T)\ :\ \kappa_{T}u\in\mathcal{P}^{r}\Lambda^{k-1}(T)\}.$ (1)
@@
The formal adjoint of $\mathrm{d}$ with respect to the $\mathrm{L}^{2}$ scalar product is denoted $\mathrm{d}^{\star}$. The Hodge star operator is denoted $\scriptstyle\bigstar$}. We have $\mathrm{d}^{\star}=\pm{\scriptstyle\bigstar}\mathrm{d}{\scriptstyle\bigstar}$.
@@
$\mathrm{d}^{\star}\mathrm{d}u=0,\quad\mathrm{d}^{\star}u=0,$ (2)
@@
Even though we adopt the FES point of view, we try to express our results first without reference to those definitions. On the other hand [21, §2] contains a glossary and a polished short presentation.
@@ 2 Lowest order case
$Z^{k}_{0}(T)$ $=\{u\in X^{k}_{0}(T)\ :\ \forall w\in W^{d-k}(T)\quad\int_{T}u\wedge w=0\}.$ (3)
@@
$0\to Z^{\scriptscriptstyle{\bullet}}_{0}(T)\to X^{\scriptscriptstyle{\bullet}}_{0}(T)\to W^{d-{\scriptscriptstyle{\bullet}}}(T)^{\star}\to 0.$ (4)
The sequence $W^{\scriptscriptstyle{\bullet}}(T)$ is exact, including at indices 0 and $d$. Hence $Z^{\scriptscriptstyle{\bullet}}_{0}(T)$ inherits the cohomology of $X^{\scriptscriptstyle{\bullet}}_{0}(T)$, by the snake lemma, in its full version [1, Theorem 3.5 page 600]. ∎
@@
$\mathrm{d}:Y^{d-1}(T)\to\ker\textstyle\int_{T}(\cdot)|_{Z^{d}(T)}.$ (7)
@@
$\int_{T}\mathrm{d}u\cdot\mathrm{d}u^{\prime}+\int_{T}\mathrm{d}u^{\prime}\wedge\kappa_{T}a+\int_{T}u^{\prime}\wedge\kappa_{T}b+\int_{T}u^{\prime}\cdot\mathrm{d}v$ $=0,$ (8)
$\int_{T}\mathrm{d}u\wedge\kappa_{T}a^{\prime}=0,\ \int_{T}u\wedge\kappa_{T}b^{\prime}$ $=0,$ (9)
$\int_{T}u\cdot\mathrm{d}v^{\prime}$ $=0,$ (10)
@@
$\mathrm{d}^{\star}\mathrm{d}u+\mathrm{d}v$ $=0,$ (11)
$\mathrm{d}^{\star}u$ $=0.$ (12)
There, discrete variants obtained from weak formulations on suitable finite dimensional spaces, were also used to make the method ”fully discrete”. The Lagrange multiplier $v$ was always 0.
@@
$\int_{T}\mathrm{d}u\cdot\mathrm{d}u^{\prime}+\int_{T}\mathrm{d}u^{\prime}\wedge\kappa_{T}a=0,$ (13)
$\int_{T}\mathrm{d}u\wedge\kappa_{T}a^{\prime}=0.$ (14)
@@
$\int_{T}u^{\prime}\wedge\kappa_{T}b+\int_{T}u^{\prime}\cdot\mathrm{d}v=0,$ (15)
$\int_{T}u\wedge\kappa_{T}b^{\prime}=0,$ (16)
$\int_{T}u\cdot\mathrm{d}v^{\prime}=0,$ (17)
@@
$\int_{T}\mathrm{d}\hat{u}\cdot\mathrm{d}u^{\prime}+\int_{T}\mathrm{d}u^{\prime}\wedge\kappa_{T}a+\int_{T}u^{\prime}\wedge\kappa_{T}b+\int_{T}u^{\prime}\cdot\mathrm{d}v$ $=-\int_{T}\mathrm{d}\tilde{u}\cdot\mathrm{d}u^{\prime},$ (18)
$\int_{T}\mathrm{d}\hat{u}\wedge\kappa_{T}a^{\prime}=-\int_{T}\mathrm{d}\tilde{u}\wedge\kappa_{T}a^{\prime},\ \int_{T}\hat{u}\wedge\kappa_{T}b^{\prime}$ $=-\int_{T}\tilde{u}\wedge\kappa_{T}b^{\prime},$ (19)
$\int_{T}\hat{u}\cdot\mathrm{d}v^{\prime}$ $=-\int_{T}\tilde{u}\cdot\mathrm{d}v^{\prime},$ (20)
@@
$\int_{T}\mathrm{d}\hat{u}\wedge\kappa_{T}a^{\prime}+\int_{T}\hat{u}\wedge\kappa_{T}b^{\prime}$ $=\int_{T}\hat{u}\wedge((-1)^{k+1}\mathrm{d}\kappa_{T}a^{\prime}+\kappa_{T}b^{\prime}),$ (21)
$=\int_{T}\hat{u}\wedge(\frac{(-1)^{k+1}}{d-k}a^{\prime}+\kappa_{T}b^{\prime}),$ (22)
@@
$w^{\prime}=\frac{(-1)^{k+1}}{d-k}a^{\prime}+\kappa_{T}b^{\prime},$ (23)
@@
$\int_{T}\mathrm{d}\hat{u}\cdot\mathrm{d}u^{\prime}+\int_{T}u^{\prime}\wedge w+\int_{T}u^{\prime}\cdot\mathrm{d}v$ $=-\int_{T}\mathrm{d}\tilde{u}\cdot\mathrm{d}u^{\prime},$ (24)
$\int_{T}\hat{u}\wedge w^{\prime}$ $=-\int_{T}\mathrm{d}\tilde{u}\wedge\kappa_{T}a^{\prime}-\int_{T}\tilde{u}\wedge\kappa_{T}b^{\prime},$ (25)
$\int_{T}\hat{u}\cdot\mathrm{d}v^{\prime}$ $=-\int_{T}\tilde{u}\cdot\mathrm{d}v^{\prime},$ (26)
for some $w\in W^{d-k}(T)$ and $v\in Y^{k-1}(T)$, and for all $w^{\prime}\in W^{d-k}(T)$ (written as (23) ) and $v^{\prime}\in Y^{k-1}(T)$.
@@
We have: $({\scriptstyle\bigstar}W^{d-k}(T))\cap\mathrm{d}Y^{k-1}(T)=0$.
@@
It holds, more generally, that for $u\in Z^{k}_{0}(T)$ we have $u\perp{\scriptstyle\bigstar}W^{d-k}(T)$ by definition. ∎

### v2

@@ Abstract
We show that the DDR method can be interpreted as defining a computable consistent discrete $\mathrm{L}^{2}$ product on a conforming FES defined by PDEs. Without modifying the numerical method itself, this point of view provides an alternative approach to the analysis. The conformity and consistency properties we obtain are stronger than those previously shown, even in low dimensions. We can also recover some of the other results that have been proved about DDR, from those that have already been proved, in principle, in the general context of FES. We also bring VEM, the Virtual Element Method, into the discussion.
@@ Introduction
We give a new presentation of the DDR method of [36, 34, 13]. We interpret it as a numerical method based on discrete spaces defined piecewise by solving certain new Partial Differential Equations (PDEs), on each cell of a mesh. The method is conforming in the sense of producing subcomplexes of natural complexes of Sobolev spaces. Even though the solutions to these new local PDEs are not computable, the numerical method itself is ”fully discrete” because it allows the explicit computation of enough moments against polynomials, from the degrees of freedom. The method is then consistent in the sense of Strang’s lemmas. This type of consistency and conformity are stronger than those previously proved for DDR. The interpretation also embeds the method in the Finite Element System framework, and allows to leverage the results already obtained in that setting.
Previous work on FES
@@
The framework of Finite Element Systems (FES) was introduced in [16] and expanded upon in [27, §5] (this independent section is available on arXiv [^1] S. H. Christiansen, Finite element systems of differential forms, arXiv:1006.4779, 2015.) and subsequent papers. It is an environment in which to define and analyze sequences of spaces of differential forms, that are finite elements in a broad sense. The underlying mesh can consist of cells (in particular polytopes) and the basis functions can be general (not necessarily based on polynomials). Thus one may call them polytopal discrete de Rham sequences, though this expression is given an a priori different meaning below. These spaces were conforming in the sense that the fields were $\mathrm{L}^{2}$-integrable with exterior derivative also $\mathrm{L}^{2}$-integrable, which enforces that traces, in the sence of pullbacks to faces, coincide (they are ”single-valued”). This encodes a type of partial continuity that is common for the mixed finite elements of electromagnetics and fluid flow [12], such as those of Raviart-Thomas-Nédélec. The latter were extended to differential forms in [43]. A synthesis with extensions to elasticity is provided in Finite Element Exterior Calculus (FEEC) [4].
@@
Since then the framework of FES has been generalized to accomodate higher regularity of the fields [25] (as in so-called ”Stokes complexes” relevant to fluid flow) and other types of fields [26] (e.g. symmetric tensors and their associated curvature tensors, as in Riemannian geometry). The definition has a natural expression in the language of categories: a FES is a just a contravariant functor from the (poset) category determined by the mesh, to the category of vector spaces (interpreted as spaces of fields on cells). A short summary is provided in [21]. This definition covers all finite elements described according to Ciarlet’s definition, including non-conforming ones (such as Morley and Crouzeix-Raviart elements).
This point of view connects finite elements with sheaf theory [46]. Indeed the combinatorial structure of the mesh can be seen as defining a ”site” with a Grothendieck topology. A contravariant functor from such a site into an algebraic category is a presheaf, which for us represents function spaces than can be both localized (from global to local) and glued back (from local to global).
@@
For the purposes of this paper, however, the definition given in [16] is sufficient. The goal back then was to define a discrete de Rham finite element sequence on polytopal meshes that mimicked the lowest order mixed finite elements that correspond to Whitney forms. Indeed these were known only for simplices (and products of simplices handled by tensor product constructions). For this presentation we will also refer to the more comprehensive presentation in [27, §5]. Short presentations can also be found in later publications, in particular [28, §2].
@@
Other polytopal methods
The Virtual Element Method (VEM), reviewed in [9], has been developed with similar goals as FES and some similar basic principles, such as definition by recursive extension of the fields, from low dimensional cells. But, crucially, no submesh is used in VEM to define the discretization spaces or the numerical method. The challenge is to reconstruct enough information on the fields from the degrees of freedom, to obtain stable and consistent approximations of the bilinear forms defining the partial differential equation to solve. A connection between the mixed VEM (defined in dimension 3) of [8] and FES (in arbitrary dimension) was made in [22, §2.1] and will be used later.
@@
The close relationship between DDR and VEM has already been pointed out in [10]. In some versions these methods can be seen to have the same degrees of freedom, and they coincide with those of FES provided in [22, §2.1].
Lowest order numerical methods have been designed on various polytopes for a long time and can be traced back to work of Yee and Arakawa in the 1960ies and 1970ies. Many seem to share a common algebraic approach to the differential operators whereas they have different approaches to the construction of the discrete $\mathrm{L}^{2}$ inner product.
Contribution
In this paper, we reinterpret DDR as a conforming FES equipped with computable consistent approximations of the $\mathrm{L}^{2}$ inner products. The degrees of freedom of [22, §2.1] are used. This shows that these examples of DDR, VEM and FES are very similar, if not identical, in practice. They seem to differ mainly in terms of interpretation and analysis. The main novelty here is just that a new harmonic extension PDE, modifying the one used in [22, §2.1], is introduced to define the discrete spaces.
This point of view allows us to recover some of the results that have been proved for VEM or DDR as special cases of results already proved, in their principle, in the general context of FES. We summarize:
@@
  • Cohomological properties developed for FES, especially [27, Proposition 5.15], now carry over to DDR. This passage from local exactness (on each cell of the mesh), to globally correct cohomology (on the computational domain) illustrates that the existence of unisolvent degrees of freedom on a FES can be interpreted as a softness condition on sheaves (see [46, §III.2] for definitions).
  • Uniformly bounded commuting projections, building on [49, 15], have been described for FEEC [4, 30] and FES [27, §5.3]. They enable one to recover Poincaré-Friedrich estimates, which are basic to the stability analysis of mixed methods.
In the setting of FES one furthermore obtains refinements in the analysis:
  • Translation estimates are given in [27, Proposition 5.71], and have applications to the equations of fluid flow.
  • Discrete Rellich compactness, as introduced by Kikuchi for Maxwell [11, §IV.19], holds generally, as proved first in [15, Corollary 5] with applications to the Hodge-Laplace eigenvalue problem on manifolds. See also [31] on the relationship of bounded commuting projections to such properties, at an abstract level.
  • Discrete Sobolev injections hold [27, Propositions 5.69, 5.70] and have applications to, for instance, non-linear Maxwell wave equations as in [29]. See also [42] on this topic.
@@
We insist that in our point of view the method is conforming in the sense that we obtain subspaces of standard Sobolev spaces and that the discrete differential operators (instances of the exterior derivative) are simply the restriction of the ordinary ones. These operators are computable in terms of the chosen degrees of freedom. Taking into account the variational crime arizing from the approximation the $\mathrm{L}^{2}$ inner product is viewed as a separate ingredient in the analysis, addressed with the help of Strang’s first lemma.
An example of a FES determined by solving local PDEs and equipped with a computable and consistent bilinear form, has been analyzed in [23]. It concerned the Schrödinger equation with a magnetic field (representing an $\mathbb{U}(1)$-connection). The local PDEs used on the cells are often related, but not identical, to the global PDE one wants to solve. This aspect can be compared with Trefftz methods which often glue exact local solutions in non-conforming ways [45].
Notice that many arguments involve a straightforward scaling combined with an appeal to compactness inspired by [4, Remark p. 64 and proof p. 69]. An optimal formulation of such compactness arguments remains to be delineated in the context of FES.
Remark.
More recent constructions of local bounded commuting projections [39, 7] carry over to FES in a very natural way. The approach of [39] extends to FES of differential forms and the approach of [7] extends to general FES as in [26] (the latter approach does not appeal to a wedge product). In fact the FES framework seems to identify exactly the algebraic properties of discretization spaces needed for these constructions. A general analysis of stability and compactness properties of these methods remains, in the context of FES, to be fleshed out. See [38] for the state of the art.
In a sense we have little to add to the design of the DDR and VEM methods themselves. Many of the defining integrations (by parts or not) against cleverly chosen polynomial forms have of course already been exemplified in the literature, though perhaps not in this order and generality. We hope the FES perspective can contribute usefully to the analysis, leading to various convergence proofs. Besides, by providing a clarified rationale for neglecting certain terms and identifying key underlying algebraic facts (such as short exact sequences of complexes and their long exact sequences of cohomology groups), this presentation can ease the extension of the method in various directions. This could be for instance the above mentioned upwinding problem, or the existing variants of DDR developed for other differential operators [37] and regularities [41].
The paper is organized as follows. In section 1 we review notations and provide a reading guide on FES. Section 2 contains the construction in the lowest order case (sequences of spaces containing constants). Section 3 contains the construction in the high order case (sequences of spaces containing polynomials of degree $r$, for any choice of $r\geq 1$). Section 4 shows how the preceding results can be used to set up a numerical method for a prototype equation and start to analyse it.
@@ 1 Notations
The subspace of $X^{k}(T)$ consisting of forms with 0 traces on $\partial T$ is denoted $X^{k}_{0}(T)$:
$X^{k}_{0}(T)=\{u\in X^{k}(T)\ :\ u|_{\partial T}=0\}.$ (1.1)
The subscript 0 will have the same meaning for other spaces as well. Elements on $X^{k}(T)$ have well defined traces, in negative index Sobolev spaces, only on codimension one faces $T^{\prime}\subseteq\partial T$. However, if these traces in turn are imposed to be in $X^{k}(T^{\prime})$, one can repeat the procedure and follow traces all the way down. One gets a space of forms with improved boundary regularity and some nice properties, such as [28, Proposition 2.4]. All our discrete spaces will have this type of extra boundary regularity.
The space of Whitney $k$-forms (which could also be attributed to Weil and de Rham) was originally presented in terms of partitions of unity, such as barycentric coordinates, but it can also be presented in terms of the Koszul operator [4, §3.2]. For each cell $T\in\mathcal{T}$, we let $b_{T}$ be its barycentre (we suppose it is ”computable”). On $T$, the Koszul operator $\kappa_{T}$ is the contraction by the vector field on $T$ given as $x\mapsto x-b_{T}$ (for $x\in T$). We adopt notations from [4]. The original lowest order Whitney forms are then $\mathcal{P}_{\!\!-}^{1}\Lambda^{k}(T)$. The high order version are denoted $\mathcal{P}_{\!\!-}^{r}\Lambda^{k}(T)$:
$\mathcal{P}_{\!\!-}^{r}\Lambda^{k}(T)=\{u\in\mathcal{P}^{r}\Lambda^{k}(T)\ :\ \kappa_{T}u\in\mathcal{P}^{r}\Lambda^{k-1}(T)\}.$ (1.2)
@@
The Hodge star operator is denoted $\scriptstyle\bigstar$. It maps from $k$-forms on $T$ to (dk)(d-k)-forms on $T$, when $d=\dim T$. It is characterized by the property that, when $u$ and $v$ are both $k$-forms:
$u\wedge\raisebox{1.0pt}{$\scriptstyle\bigstar$}v=(u\cdot v)\,\omega,$ (1.3)
where the pointwise scalar product is denoted $(\cdot)$ and the volume form on $T$ associated with the scalar product and the orientation is denoted $\omega$. For instance $\raisebox{1.0pt}{$\scriptstyle\bigstar$}1=\omega$. The formal adjoint of $\mathrm{d}$ with respect to the $\mathrm{L}^{2}$ scalar product is denoted $\mathrm{d}^{\star}$. We have $\mathrm{d}^{\star}=\pm\raisebox{1.0pt}{$\scriptstyle\bigstar$}\mathrm{d}\raisebox{1.0pt}{$\scriptstyle\bigstar$}$.
@@
$\mathrm{d}^{\star}\mathrm{d}u=0,\quad\mathrm{d}^{\star}u=0,$ (1.4)
@@
Even though we adopt the FES point of view, we first present our results without relying on those definitions. On the other hand [28, §2] contains a glossary and a short presentation that is more than sufficient for this article.
FES: an overview
Since the FES framework has not been reviewed recently we now offer a guide to the relevant literature, for the convenience it might provide the reader.
  • In [16] the FES framework was introduced as auxiliary discrete spaces inside which analogues of (lowest order) Whitney forms could be defined as locally harmonic forms.
  • Reference [17] is mostly expository. In §3.1 some background in algebra is provided (the five lemma, the snake lemma and the long exact cohomology sequence, the Künneth theorem on tensor products). In §3.2 basic concepts in differential geometry are reviewed. This algebraic and differential geometric background is then applied to study FES. The motivation was the construction of Galerkin methods for Maxwell’s equations, complementing [44, 47, 48] on some points, particularly in regards to general meshes and basis functions.
    A ”geometric decomposition” in terms of ”bubble spaces”, in the sense of [5], is given in §3.3.2 (Proposition 3.14) as a byproduct of discretely harmonic forms.
    A first study of tensor products of FES is provided in §3.3.3 and further developed in later publications. This theory covers standard elements as in [3, §3] and splines as in [14]. In fact, one can start with any space of functions on an interval that contains a non-trivial constant function and one with different values at the endpoints. The latter function can be for instance an exponential.
  • In [27, §5] the framework was further developed in its own right, to account for the mixed finite element spaces synthesized in FEEC. In §5.3 commuting quasi-interpolators deduced from smoothings were studied. In §5.4 stability properties were deduced.
  • In [19] it is illustrated that by modifying the definition of harmonicity one can get upwinded complexes of differential forms, useful for convection diffusion problems. The above mentioned exponential function allows one to recover some tensor product “exponentially fitted” methods. The scalar convection diffusion problem is analyzed in [24] in fractional Sobolev spaces (of the type $\mathrm{H}^{1/2}$).
  • In [28] FES was used to describe the spaces $\mathcal{P}_{\!\!-}^{r}\Lambda^{k}(T)$ of trimmed polynomial differential forms on simplices. In §2 a glossary and a seven page summary of FES is provided. Alternative degrees of freedom were proposed. Resolutions were introduced to describe relations in the natural spanning families constructed with barycentric coordinates.
    It is believed that working with canonical generators and relations (instead of selecting a specific basis) and systematically treating restrictions to subcells, could be valuable for implementations, not just the theoretical study.
  • In [22] it was shown that the Serendipity element [2], the TNT element [33] and the trimmed polynomial differential forms are all examples of FES satisfying a minimality condition. A connection with VEM was made explicit in §2.1. A general recipy for constructing minimal or ”serendipity” spaces was given in §3.
  • In [25] and [26] the framework of FES is generalized to account for more types of finite element spaces. In [26, 21] the connection with category theory and sheaves is described.
@@ 2 Lowest order case
$Z^{k}_{0}(T)$ $=\{u\in X^{k}_{0}(T)\ :\ \forall w\in W^{d-k}(T)\quad\int_{T}u\wedge w=0\}.$ (2.1)
@@
$0\to Z^{\scriptscriptstyle{\bullet}}_{0}(T)\to X^{\scriptscriptstyle{\bullet}}_{0}(T)\to W^{d-{\scriptscriptstyle{\bullet}}}(T)^{\star}\to 0.$ (2.2)
The sequence $W^{\scriptscriptstyle{\bullet}}(T)$ is exact, including at indices 0 and $d$. Hence $Z^{\scriptscriptstyle{\bullet}}_{0}(T)$ inherits the cohomology of $X^{\scriptscriptstyle{\bullet}}_{0}(T)$, by the snake lemma, in its full version. See [1, Theorem 3.5, p. 600] or [17, Theorem 3.1, p. 356]. ∎
@@
$\mathrm{d}:Y^{d-1}(T)\to\ker\textstyle\int_{T}(\cdot)|_{Z^{d}(T)}.$ (2.5)
@@
$\int_{T}\mathrm{d}u\cdot\mathrm{d}u^{\prime}+\int_{T}\mathrm{d}u^{\prime}\wedge\kappa_{T}a+\int_{T}u^{\prime}\wedge\kappa_{T}b+\int_{T}u^{\prime}\cdot\mathrm{d}v$ $=0,$ (2.6)
$\int_{T}\mathrm{d}u\wedge\kappa_{T}a^{\prime}=0,\ \int_{T}u\wedge\kappa_{T}b^{\prime}$ $=0,$ (2.7)
$\int_{T}u\cdot\mathrm{d}v^{\prime}$ $=0,$ (2.8)
@@
$\mathrm{d}^{\star}\mathrm{d}u+\mathrm{d}v$ $=0,$ (2.9)
$\mathrm{d}^{\star}u$ $=0.$ (2.10)
The Lagrange multiplier $v$ was always 0. This can be interpreted as solving a Hodge-Laplace problem. Indeed, with $\Delta=\mathrm{d}^{\star}\mathrm{d}+\mathrm{d}\mathrm{d}^{\star}$, the above system is equivalent to:
$\Delta u=0,\ \textrm{and }\ (\mathrm{d}^{\star}u)|_{\partial T}=0.$ (2.11)
In both cases the boundary condition is that $u|_{\partial T}$ is given (in the sense of pullbacks).
In [16], discrete variants obtained from weak formulations on suitable finite dimensional spaces, were also used to make the method ”fully discrete”.
To sum up the definition of $B^{k}(T)$ involves a variant of the Hodge-Laplace equation deduced for certain boundary conditions, modified by some polynomial constraints.
@@
$\int_{T}\mathrm{d}u\cdot\mathrm{d}u^{\prime}+\int_{T}\mathrm{d}u^{\prime}\wedge\kappa_{T}a=0,$ (2.12)
$\int_{T}\mathrm{d}u\wedge\kappa_{T}a^{\prime}=0.$ (2.13)
@@
$\int_{T}u^{\prime}\wedge\kappa_{T}b+\int_{T}u^{\prime}\cdot\mathrm{d}v=0,$ (2.14)
$\int_{T}u\wedge\kappa_{T}b^{\prime}=0,$ (2.15)
$\int_{T}u\cdot\mathrm{d}v^{\prime}=0,$ (2.16)
@@
$\int_{T}\mathrm{d}\hat{u}\cdot\mathrm{d}u^{\prime}+\int_{T}\mathrm{d}u^{\prime}\wedge\kappa_{T}a+\int_{T}u^{\prime}\wedge\kappa_{T}b+\int_{T}u^{\prime}\cdot\mathrm{d}v$ $=-\int_{T}\mathrm{d}\tilde{u}\cdot\mathrm{d}u^{\prime},$ (2.17)
$\int_{T}\mathrm{d}\hat{u}\wedge\kappa_{T}a^{\prime}=-\int_{T}\mathrm{d}\tilde{u}\wedge\kappa_{T}a^{\prime},\ \int_{T}\hat{u}\wedge\kappa_{T}b^{\prime}$ $=-\int_{T}\tilde{u}\wedge\kappa_{T}b^{\prime},$ (2.18)
$\int_{T}\hat{u}\cdot\mathrm{d}v^{\prime}$ $=-\int_{T}\tilde{u}\cdot\mathrm{d}v^{\prime},$ (2.19)
@@
$\int_{T}\mathrm{d}\hat{u}\wedge\kappa_{T}a^{\prime}+\int_{T}\hat{u}\wedge\kappa_{T}b^{\prime}$ $=\int_{T}\hat{u}\wedge((-1)^{k+1}\mathrm{d}\kappa_{T}a^{\prime}+\kappa_{T}b^{\prime}),$ (2.20)
$=\int_{T}\hat{u}\wedge(\frac{(-1)^{k+1}}{d-k}a^{\prime}+\kappa_{T}b^{\prime}),$ (2.21)
@@
$w^{\prime}=\frac{(-1)^{k+1}}{d-k}a^{\prime}+\kappa_{T}b^{\prime},$ (2.22)
@@
$\int_{T}\mathrm{d}\hat{u}\cdot\mathrm{d}u^{\prime}+\int_{T}u^{\prime}\wedge w+\int_{T}u^{\prime}\cdot\mathrm{d}v$ $=-\int_{T}\mathrm{d}\tilde{u}\cdot\mathrm{d}u^{\prime},$ (2.23)
$\int_{T}\hat{u}\wedge w^{\prime}$ $=-\int_{T}\mathrm{d}\tilde{u}\wedge\kappa_{T}a^{\prime}-\int_{T}\tilde{u}\wedge\kappa_{T}b^{\prime},$ (2.24)
$\int_{T}\hat{u}\cdot\mathrm{d}v^{\prime}$ $=-\int_{T}\tilde{u}\cdot\mathrm{d}v^{\prime},$ (2.25)
for some $w\in W^{d-k}(T)$ and $v\in Y^{k-1}(T)$, and for all $w^{\prime}\in W^{d-k}(T)$ (written as (2.22) ) and $v^{\prime}\in Y^{k-1}(T)$.
@@
We have: $(\raisebox{1.0pt}{$\scriptstyle\bigstar$}W^{d-k}(T))\cap\mathrm{d}Y^{k-1}(T)=0$.
@@
It holds, more generally, that for $u\in Z^{k}_{0}(T)$ we have $u\perp\raisebox{1.0pt}{$\scriptstyle\bigstar$}W^{d-k}(T)$ by definition. ∎